# AI Flow: Perspectives, Scenarios, and Approaches

### v2

@@ I-A The Rise of AI
The concept of AI, first founded as an academic discipline in 1956, has undergone transformative growth through iterative advancements. Early progress in AI centered on foundational machine learning (ML) theories, including regression analysis, support vector machines (SVMs), and artificial neural networks. Driven by the rapid development in computational power and data availability, breakthroughs in deep learning later revolutionized the field. Pioneering models like AlexNet and ResNet attained unprecedented accuracies in image classification, object detection, and speech recognition, catalyzing their widespread adoption across industries such as healthcare, robotics, and autonomous systems. Reinforcement learning (RL) demonstrated its superhuman capabilities in domains requiring long-term planning as well.
@@ III Familial Model
The familial model refers to a series of different-sized models with aligned hidden features, thereby enabling overhead-free information sharing and effective collaboration. This architecture supports an almost arbitrary number of parameters, allowing it to fully exploit the computational capabilities of heterogeneous devices and satisfy the demands of various downstream tasks. Contrary to conventional models of varying sizes, familial models can directly reuse the intermediate results from smaller models without additional middleware. Consequently, the cooperative inference of multiple familial models exhibits remarkable efficiency, particularly under stringent constraints on communication bandwidth and hardware resources. In the following subsections, we first elaborate on two enabling techniques of familial models: weight decomposition and early exit. We then propose two strategies to implement familial models for LLMs and VLMs based on the aforementioned enabling techniques. Experimental results on trained familial models are presented to demonstrate their performance and adaptability. Finally, discussions on the benefits and applications of familial models are given.
III-A Enabling Techniques
III-A1 Weight Decomposition
@@
$L$ $=\left\|\bm{W}\bm{X}-\bm{W}_{\rm u}\bm{W}_{\rm v}\bm{X}\right\|_{F}^{2}=\left\|\sum_{i=h+1}^{r}\sigma_{i}\bm{u}_{i}\bm{v}_{i}^{T}\bm{S}^{-1}\bm{X}\right\|_{F}^{2}=\sum_{i=h+1}^{r}\sigma_{i}^{2},$ (3)
@@
III-A2 Early Exit
The second enabling technique of familial models is early exit [38, 39], where intermediate features produced by earlier layers are directly forwarded to the language model (LM) head to predict the output token. This design enables LLM inference on resource-constrained edge devices without significantly compromising accuracy. The LM head consists of a layer normalization module and a linear layer that transforms the intermediate features into probabilities of all tokens in the vocabulary. To further improve model performance at earlier exit points, we introduce lightweight trainable modules between each exit point and the LM head to refine the intermediate features before they are used for prediction, referred to as branches. Multiple branches with varying numbers of parameters and computational complexities can be connected to a single exit point, enabling flexible adaptation to diverse hardware capabilities and task requirements. For instance, transformer blocks whose linear layers are decomposed using different compression ratios can be inserted between the exit point and the LM head, forming a series of branches at each exit point. Consequently, the familial model supports almost arbitrary model capabilities while maintaining consistency across different configurations. Heterogeneous devices can adaptively choose the optimal branch based on the available hardware resources and the delay requirements of downstream tasks.
@@ III-A2 Early Exit
III-B Implementation Strategies
III-B1 Hierarchical Principal Component Decomposition
@@
Details. We represent the weight matrix in each linear layer as a weighted sum of $K$ low-rank components. This decomposition can be interpreted as a matrix-valued Taylor-like series expansion, where each low-rank term progressively approximates the residual error left by the previous components. The HPCD training framework utilizes a hierarchical initialization strategy followed by iterative residual fitting. This design enables each component to sequentially capture critical features while preserving the information learned by the previous components. For the initialization of the first low-rank component, we employ the weight decomposition method described in Section III-A1, which decomposes the linear layer weight matrices of a pre-trained model. For subsequent components, we adopt a sequential residual approximation approach. At each step $k\geq 2$, the parameters of all previously trained components are frozen, and the residual (defined as the difference between the original weight matrix and the cumulative approximation of the prior components) is computed. To initialize the $k$-th component, this residual matrix is approximated using the aforementioned low-rank decomposition method. This iterative refinement process ensures that each new component enhances the overall approximation by maintaining orthogonality with respect to earlier components. During inference, a compact weight matrix is obtained through HPCD by retaining the first-$k$ components. The value of $k$ is adjusted based on compute or hardware constraints, allowing granular control over FLOPs and memory usage. We design four model variants with parameter scales ranging from 2.38B to 6.30B. All variants share a uniform low-rank decomposition structure where each component maintains a consistent rank, while the number of components $k$ varies to achieve different parametric capacities. Specifically, we adapt the TeleChat [40] architecture by substituting the weight matrices in self-attention and feed-forward layers with HPCD-structured components. For models with parameter sizes ranging from 2.38B to 6.30B, we select the number of components $k$ with values ranging from $1$ to $4$.
@@
We employ a large-scale, high-quality pre-training dataset curated through aggregation from diverse sources, establishing a comprehensive knowledge base. To preserve contextual coherence in extended sequences, we concatenate documents from the same source. Sequences are truncated and concatenated to achieve a fixed maximum context length of 8K tokens. In terms of training methodology, we utilize the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$. For learning rate scheduling, we implement a cosine decay schedule with a peak learning rate of $3\times 10^{-4}$, which linearly decays to 10% of the peak value by the end of training. To improve training stability, we apply global gradient norm clipping of 1.0. Additionally, a weight decay of 0.01 is applied to all trainable parameters (excluding biases) to regularize the model. We employ a batch size of 4M tokens during training, with the total training token counts (in billions) ranging from $400$ to $100$.
@@
III-B2 Early Exiting with Scalable Branches
@@
We propose the second implementation strategy of familial models, i.e., early exiting with scalable branches (EESB), to minimize the design cost of familial models while supporting a nearly arbitrary number of parameters. As illustrated in Fig. 8, this strategy inserts a branch network between each exit point and the shared LM head. The branch network is formulated as a decomposed transformer block, whose linear layers are decomposed as elaborated in Section III-A1, and the total parameter count is equal to that of an original transformer block. Based on the parameter budget, the linear layers of the transformer block in the branch network can be further decomposed with a lower compression ratio, achieved by retaining only a portion of the decomposed weight matrices. This design enables a nearly arbitrary number of parameters while maintaining the total parameter count of a series of familial models below twice that of the main model.
@@
III-C Discussions
The advantages of familial models are twofold: highly flexible parameter configuration and enhanced efficiency in model collaboration. First, familial models support fine-grained adjustments on the total parameter count, attained by adaptive weight decomposition of linear layers within transformer blocks. By carefully tuning the dimensions of hidden features between decomposed layers, a nearly arbitrary number of parameters is achievable for familial models, accommodating the hardware capabilities of heterogeneous devices. Given a giant model with tens of billions parameters that runs on powerful computing clusters in the cloud infrastructure, much smaller familial models can be customized to enable deployment on resource-constrained user devices and edge servers. Apart from adapting to available computing resources, familial models are also adaptive to the diverse requirements of downstream tasks on both model intelligence and inference latency. For instance, for delay-sensitive tasks such as simultaneous interpretation and autonomous robotic control, employing a smaller model can reduce response latency and enhance user experience, which is critical to model adoption in real-world scenarios.

### v3

@@ Abstract
Keywords
AI Flow, distributed inference, edge artificial intelligence, large language model, collective intelligence
@@ I-A The Rise of AI
The concept of AI, first founded as an academic discipline in 1956, has undergone transformative growth through iterative advancements. Early progress in AI centered on foundational machine learning (ML) theories, including regression analysis, support vector machines (SVMs), and artificial neural networks. Driven by the rapid development in computational power and data availability, breakthroughs in deep learning later revolutionized the field [11, 12, 13, 14, 15]. Pioneering models like AlexNet and ResNet attained unprecedented accuracies in image classification, object detection, and speech recognition, catalyzing their widespread adoption across industries such as healthcare, robotics, and autonomous systems. Reinforcement learning (RL) demonstrated its superhuman capabilities in domains requiring long-term planning as well.
@@ III Familial Model
The familial model refers to a series of different-sized models with aligned hidden features, thereby enabling overhead-free information sharing and effective collaboration. This architecture supports an almost arbitrary number of parameters, allowing it to fully exploit the computational capabilities of heterogeneous devices and satisfy the demands of various downstream tasks. Contrary to conventional models of varying sizes, familial models can directly reuse the intermediate results from smaller models without additional middleware. Consequently, the cooperative inference of multiple familial models exhibits remarkable efficiency, particularly under stringent constraints on communication bandwidth and hardware resources.
In the following subsections, we first introduce related works and then elaborate on two enabling techniques of familial models: weight decomposition and early exit. We propose two strategies to implement familial models for LLMs and VLMs based on the aforementioned enabling techniques. Experimental results on familial models are presented to demonstrate their performance and adaptability. Finally, discussions on the benefits and applications of familial models are given.
III-A Related Works
III-A1 Weight Decomposition for Parameter Efficiency
Low-rank decomposition is widely used to reduce model parameters while maintaining structural consistency. Early work by Sainath et al. [42] applied low-rank matrix factorization to speech models, while Yang et al. [43] explored training low-rank structures directly via SVD to preserve performance. In this work, we develop upon these SVD-based approaches by introducing data whitening and truncating the components with the smallest singular values (Section III-B1), enabling dynamic control of the decomposition rank per layer. This design ensures fine-grained parameter tuning while minimizing the distortion caused by approximating the original weight matrices. Another line of approaches to enhance parameter efficiency by weight decomposition is LoRA, first proposed by Hu et al. [35], which has become a widely adopted technique for efficiently fine-tuning large pre-trained models. It freezes the backbone parameters and inserts low-rank matrices for adaptation, achieving performance parity with significantly fewer trainable parameters. Recent improvements, such as LoRA+ by Hayou et al. [44], optimize learning rates and scaling for even larger models. In contrast to modern LoRA applications, our familial model architecture extends the concept to scalable branch networks, combining low-rank decomposed transformer blocks with shared LM heads to support efficient multi-branch inference and fine-grained tuning on the parameter count.
III-A2 Early-Exit Methods
AI models typically make predictions by processing the input data through a series of structurally similar layers and exiting at the last layer. To speed up inference and reduce computational costs, many studies [45, 46, 47] incorporate multiple exit points with side branches, allowing early exiting at an intermediate layer to reduce computational complexity. Specifically, BranchyNet by Teerapittayanon et al. [48] pioneered this idea by adding side branches to early layers. Branchy-GNN proposed by Shao et al. [49] extended this approach to efficient point cloud processing based on graph neural networks. Liu et al. [50] proposed the first unified and end-to-end approach for anytime prediction, which attaches a cascade of exit points to make predictions from intermediate features. More recent studies on employing early exit in transformer-based models include DeeBERT [51], SkipBERT [52], and CALM [53]. Notably, EE-LLM [46] implemented large-scale early-exit language models with 3D parallelism and KV cache compatibility. Pan et al. introduced a lightweight and economical solution to fine-tune early-exit LLMs [54]. In this work, we propose a novel framework that introduces model decomposition to the branch networks after exit points. This design enables flexible parameter control while preserving alignment across branches, thereby supporting collaborative inference on heterogeneous devices.
III-B Enabling Techniques
III-B1 Weight Decomposition
@@
$L_{\rm decomp}$ $=\left\|\bm{W}\bm{X}-\bm{W}_{\rm u}\bm{W}_{\rm v}\bm{X}\right\|_{F}^{2}=\left\|\sum_{i=h+1}^{r}\sigma_{i}\bm{u}_{i}\bm{v}_{i}^{T}\bm{S}^{-1}\bm{X}\right\|_{F}^{2}=\sum_{i=h+1}^{r}\sigma_{i}^{2},$ (3)
@@
III-B2 Early Exit
The second enabling technique of familial models is early exit [56, 49, 45], where intermediate features produced by earlier layers are directly forwarded to the language model (LM) head to predict the output token. This design enables LLM inference on resource-constrained edge devices without significantly compromising accuracy. The LM head consists of a layer normalization module and a linear layer that transforms the intermediate features into probabilities of all tokens in the vocabulary. To further improve model performance at earlier exit points, we introduce lightweight trainable modules between each exit point and the LM head to refine the intermediate features before they are used for prediction, referred to as branches. Multiple branches with varying numbers of parameters and computational complexities can be connected to a single exit point, enabling flexible adaptation to diverse hardware capabilities and task requirements. For instance, transformer blocks whose linear layers are decomposed using different compression ratios can be inserted between the exit point and the LM head, forming a series of branches at each exit point. Consequently, the familial model supports almost arbitrary model capabilities while maintaining consistency across different configurations. Heterogeneous devices can adaptively choose the optimal branch based on the available hardware resources and the delay requirements of downstream tasks.
@@ III-B2 Early Exit
III-C Implementation Strategies
III-C1 Hierarchical Principal Component Decomposition
@@
Details. We represent the weight matrix in each linear layer as a weighted sum of $K$ low-rank components. This decomposition can be interpreted as a matrix-valued Taylor-like series expansion, where each low-rank term progressively approximates the residual error left by the previous components. The HPCD training framework utilizes a hierarchical initialization strategy followed by iterative residual fitting. This design enables each component to sequentially capture critical features while preserving the information learned by the previous components. For the initialization of the first low-rank component, we employ the weight decomposition method described in Section III-B1, which decomposes the linear layer weight matrices of a pre-trained model. For subsequent components, we adopt a sequential residual approximation approach. At each step $k\geq 2$, the parameters of all previously trained components are frozen, and the residual (defined as the difference between the original weight matrix and the cumulative approximation of the prior components) is computed. To initialize the $k$-th component, this residual matrix is approximated using the aforementioned low-rank decomposition method. This iterative refinement process ensures that each new component enhances the overall approximation by maintaining orthogonality with respect to earlier components. During inference, a compact weight matrix is obtained through HPCD by retaining the first-$k$ components. The value of $k$ is adjusted based on compute or hardware constraints, allowing granular control over FLOPs and memory usage. We design four model variants with parameter scales ranging from 2.38B to 6.30B. All variants share a uniform low-rank decomposition structure where each component maintains a consistent rank, while the number of components $k$ varies to achieve different parametric capacities. Specifically, we adapt the TeleChat [57] architecture by substituting the weight matrices in self-attention and feed-forward layers with HPCD-structured components. For models with parameter sizes ranging from 2.38B to 6.30B, we select the number of components $k$ with values ranging from $1$ to $4$.
@@
We employ a large-scale, high-quality pre-training dataset curated through aggregation from diverse sources, establishing a comprehensive knowledge base. To preserve contextual coherence in extended sequences, we concatenate documents from the same source. Sequences are truncated and concatenated to achieve a fixed maximum context length of 8K tokens. In terms of training methodology, we utilize the Adam optimizer with $\beta_{1}=0.9$, and $\beta_{2}=0.95$. For learning rate scheduling, we implement a cosine decay schedule with a peak learning rate of $3\times 10^{-4}$, which linearly decays to 10% of the peak value by the end of training. To improve training stability, we apply global gradient norm clipping of 1.0. Additionally, a weight decay of 0.01 is applied to all trainable parameters (excluding biases) to regularize the model. We employ a batch size of 4M tokens during training, with the total training token counts (in billions) ranging from $400$ to $100$.
@@
III-C2 Early Exiting with Scalable Branches
@@
We propose the second implementation strategy of familial models, i.e., early exiting with scalable branches (EESB), to minimize the design cost of familial models while supporting a nearly arbitrary number of parameters. As illustrated in Fig. 8, this strategy inserts a branch network between each exit point and the shared LM head. The branch network is formulated as a decomposed transformer block, whose linear layers are decomposed as elaborated in Section III-B1, and the total parameter count is equal to that of an original transformer block. Based on the parameter budget, the linear layers of the transformer block in the branch network can be further decomposed with a lower compression ratio, achieved by retaining only a portion of the decomposed weight matrices. This design enables a nearly arbitrary number of parameters while maintaining the total parameter count of a series of familial models below twice that of the main model. Compared with the previous study EE-LLM [46], the advantage of EESB is the ability to scale the parameter budget more flexibly. In addition to multiple exit points, each branch network contains a decomposed transformer block, where its parameters can be dynamically tuned to achieve a more flexible tradeoff between computational cost and model performance, based on task requirements or resource constraints.
@@
III-D Discussions
Guided by the valuable insights above, the familial models exhibit two advantages: highly flexible parameter configuration and enhanced efficiency in model collaboration. First, familial models support fine-grained adjustments on the total parameter count, attained by adaptive weight decomposition of linear layers within transformer blocks. By carefully tuning the dimensions of hidden features between decomposed layers, a nearly arbitrary number of parameters is achievable for familial models, accommodating the hardware capabilities of heterogeneous devices. Given a giant model with tens of billions parameters that runs on powerful computing clusters in the cloud infrastructure, much smaller familial models can be customized to enable deployment on resource-constrained user devices and edge servers. Apart from adapting to available computing resources, familial models are also adaptive to the diverse requirements of downstream tasks on both model intelligence and inference latency. For instance, for delay-sensitive tasks such as simultaneous interpretation and autonomous robotic control, employing a smaller model can reduce response latency and enhance user experience, which is critical to model adoption in real-world scenarios.
@@ VII Conclusions
Abbreviations
Abbreviation Definition
AI Artificial Intelligence
IT Information Technology
CT Communication Technology
LLM Large Language Model
IoT Internet of Thing
VLM Vision-Language Model
ML Machine Learning
SVM Support Vector Machine
RL Reinforcement Learning
RNN Recurrent Neural Network
CNN Convolutional Neural Network
CoT Chain-of-Thought
MAS Multi-Agent System
GB Gigabyte
BS Base Station
RSU Roadside Unit
TOFC Task-Oriented Feature Compression
DPC-KNN Density Peaks Clustering Based on $K$ Nearest Neighbors
PDF Probability Density Function
STE Straight-Through Estimator
MLP Multi-Layer Perceptron
FFN Feed-Forward Network
SVD Singular Value Decomposition
LM Language Model
HPCD Hierarchical Principal Component Decomposition
KL Kullback-Leibler
EESB Early Exiting with Scalable Branches
VQA Visual Question Answering
OCR Optical Character Recognition
INS Interleaved Interaction Synthesis
REC Relative Coordination Refinement
FID Fréchet Inception Distance
FVD Fréchet Video Distance
VAE Variational Autoencoder
KVD Kernel Video Distance
GUI Graphical User Interface
AR Augmented Reality
FL Federated Learning
PEFT Parameter-Efficient Fine-Tuning
WANET Wireless Ad Hoc Networks
TABLE X: List of abbreviations and their definitions.
Table X is a list of abbreviations used throughout this paper.
Declarations
Availability of data and material
While most of the datasets used for training are publicly available, a portion of training data for HPCD models originates from proprietary datasets and cannot be shared due to legal and contractual restrictions. These restricted datasets were used exclusively for model training and do not affect the reproducibility of the core findings presented in this paper. Source data are provided with this paper.
Code availability
Code for the implementation of familial models is publicly available on GitHub at https://github.com/TeleAI-AI-Flow/AI-Flow-Ruyi.
Competing interests
The authors declare no competing interests.
Funding
The authors declare no applicable fundings.
Authors’ contributions
Xuelong Li initiated and led this research project. Jiawei Shao facilitated this work. Hongjun An, Wenhan Hu, Sida Huang, Siqi Huang, Ruanjun Li, Yiliang Song, Zihan Wang, and Cheng Yuan contributed to the experiments with the assistance from Jiawei Shao, Hongyuan Zhang, and Yuanzhi Liang, under the guidance of Chi Zhang. Wenhao Zhuang, Ruanjun Li, Yuanzhi Liang, Cheng Yuan, Siqi Huang, Zihan Wang, and Jiawei Shao wrote this paper.
Acknowledgements
We thank Yunuo Hu for assistance in the illustration of some figures, and Chen Chen for assistance in the revision of some references.